Ukted Odwolanis Korespondenga RecencaWidok Pomoc P Powiedz mi co chcesz zrobi Fuedamentals of Quantitative Methods Part IV (Exercises from Business Sharistics: Sharpe, De Veaux, Velleman) 1. You and your friend decide to get your ears inspected. You are informed that 79% of cars pass inspection. If the event of your cars passing is indepeodent of your frieeds cr a) What is the probability that your car passes inspection? b) What is the probability that your car doesnt pass inspection? c) What is the probabelity that both of the cars pass? d) What is the probability thst a least one of the two cars passes Quality control. A tire manufacturer recently announced a recall because 2% of its tires are defective. If you just bought a new set of four tires from this manufactarer, what is the probability that at least ooe of your new tires is defective? 2. Pepsi promotion. For a sales promotion, the manufacturer places witing symbols under the caps of 10% of all Pepsi bottles. If you buy a sjx-pack of Pepsi, what is the probability that you win soanething? 3. 4. Auto warrasty, In developing their warracty policy, an atomobile coumpany estimates that over a I-yesr period 17% of their tiew cars will meed to be repaired once. 7% will need repairs trice, and 4% will require three or more repairs If you bry s new car from them, what is the probability that your car will need a) No repairs? bi No more than oue repair? c) Some repaies If you boughr two new cars, what is the probability that: a) Neither will need repair? b) Both will need repair? e) At least one car will weed repair? 5. Consulting team You work for a large global management cousulting compazy. Of the entire work force 55% lave had no experience in the teleconumunicatious industry, 32% lave had limited of analysts, experience (less than 5 years), and the rest have had extensive experience (5 years or more). On a receut project. you an the peoject involves twlecommicaticns. What is the probability that the first teanumate you meet has a) Extensive telecomenunications experience? by Some telecommnications experience?c) No move th lamited teleconmunications experience? What is the probability that of your other two tenm mates a) Neitber has any telecomzonications experience? bi Both have some telecomusosications experience c) At least one lhas bad estensive telecoamunications experience? d two other analysts were cbosen at raudom to coustitute a team. It tonss out that part of

EA 18 FOMClass4- widok chroniony-Zapisano w: ten komputer Układ Odwołania Korespondencja Recenzja Widok Pomoc ρ Powiedz mi, co chcesz zrobić internetowe moze zawierać wirusy. Jeśli nie ma konieczności jego edytowania bezpieczniej est pozostać w widoku chronionym wiat expenence (less than 5 years), and the rest have had extensive experience (5 years or more). On a recent project, you and two other analysts were chosen at random to constitute a team. It turns out that part of What is the probability that the first teammate you meet has: e project involves a) Extensive telecommunications experience? b) Some telecommunications experience? c) No more than limited telecommunications experience? What is the probability that of your other two team mates: a) Neither has any telecommunications experience? b) Both have some telecommunications experience? c) At least one has had extensive telecommunications experience? Human resource data. Employment data at a large company reveal that 72% of the workers are married, 44% are college graduates, and half of the college grads are married. Whats the probability that a randomly chosen worker is: a) Neither married nor a college graduate? b) Married but not a college graduate? c) Married or a college graduate? 6. Mars product information. The Mars company says that before the introduction of purple, yellow made up 20% of their plain M&M candies, red made up another 20%, and orange, blue, and green each made up 10%. The rest were brown. a) If you picked an M&M at random from a pre-purple bag of candies, what is the probability that it was: i) Brown? ii) Yellow or orange? ii) Not green? iv) Striped 7. b) Assuming you had an infinite supply of M&Ms with the older color distribution, if you picked three M&Ms in a row, what is the probability that i) They are all brown? i) The third one is the first one thats red? ii) None are yellow? iv) At least one is green? c) If you draw one M&M, are the events of getting a red one and getting an orange one disjoint or independent or neither? d) If you draw two M&Ms one after the other, are the events of getting a red on the first and a red on the second disjoint or independent or neither? American Red Cross. The American Red Cross must track their supply and demand for various blood types They estimate that about 45% of the US, population has Type O blood, 40, Type A. I 1% Type 8.

าธ FOMCass4- widok chroniony-Zapisano w ten Rompuner Odwotaria Korespondencja Recenzja widok Pomoc ρ Powedz mi, co chcese sobie 8. American Red Cross. The American Red Cross nust track their supply and demand for various blood types. They estimate that about 45% of the US. population bas Type O blood, 40% Type A, l 1% Type B, and the rest Type AB a) If someone volunteers to give blood, what is the probabality that this donor i) Has Type AB blood? ii) Has Type A er Type B blood? iii) Is not Type O? b) Among four podential donors, what is the probability that: i) All are Type O? ii) None have Type AB bleed? isi) Not all are Type A7 iv) At least one person is Type B7 c) If you examine one donor, are the events of the donor being Type A and the donor being Type B disjoint or independeet or neitber? Explain your answer d) If you examine two donors, are the events that the first donor is Type A and the secood donor is Type B disjoint ce independent or oeither? 9 An orthodontist bas three fincing paclages, and each has a different service charge. He estimates that 30% of patients use the first plan, which has S10 finance charge, 50% use the second plan, which has a 520 finance charge, and 20% use the third plan, which has a S 30 finance charge. a) Find the expected vahue of the service charge b) Find the standard deviation of the service charge 10 Given indepeadent random variables. X and Y mwith means Ex-10. EY-20 and standard deviations sD(Xy-2-SDY)-S Find the mean and standard deviation of each of the variable 11 A broker has calculated the expected values of two different finacial instruments X and suppose that EX-s100. EY-90, SDN-12and SDY-$8. Find each of the following a) Eix-10) and SDIX+101 bi EiSY) and SDISY c) E(2X-Y) aud SDX2X-Y d) What assumption amust you male in part e? 12. Lottery. Jona has a lottery game called Pick 3 in which castoeners buy a ticket for SI aud choose turee munbers, each from zero to nine. They also mast select the play type, which deternines what combicatioes ave winners. In one type of play, called the Straight Bos they wi if they match the thwee eambers in auy order, but the payout is greater if the order is exact For the case where all hree of the cumbers selected are datherent, the probabalities and payoats ate Probabiity Payoar Straight Bx in 1ooo $330

FOMClass4 widok chroniony - Zapisano w. ten komputer Układ Odwołania Korespondencja Recenzja Widok Pomoc O Powiedz mi, co chcesz zrobić temetowej i moze zawierać wirusy. leśli nie ma koniecznošici jego edytowania, bezpieczniej jest pozostać wwidoku chronionym. Wącz ed 11. A broker has calculated the expected values of two different financial instruments X and Y. Suppose that ECX)-$100, E(Y)-$90, SD(X) $12 and SD(Y)-$8. Find each of the following: a) E(X+10) and SD(X+10); b) E(5Y) and SD(5Y) c) E(2X+ Y) and SD(2X+Y); d) What assumption must you make in part c)? 12. Lottery. Iowa has a lottery game called Pick 3 in which customers buy a ticket for $1 and choose three numbers, each from zero to nine. They also must select the play type, which determines what combinations are winners. In one type of play, called the Straight/Box, they win if they match the three numbers in any order, but the payout is greater if the order is exact. For the case where all three of the numbers selected are different, the probabilities and payouts are: Probability Payout Straight/Boxin 1000$350 Exact Straight/Box5 in 1000 $50 Any a) Find the amount a Straight/Box player can expect to win. b) Find the standard deviation of the players winnings. c) Tickets to play this game cost $1 each. If you subtract SI from the result in part a), what is the expected result of playing this game? 13. Bike sale. A bicycle shop plans to offer 2 specially priced childrens models at a sidewalk sale. The basic model will return a profit of S120 and the deluxe model S150. Past experience indicates that sales of the basic model will have a mean of 5.4 bikes with a standard deviation of 1.2, and sales of the deluxe model will have a mean of 3.2 bikes with a standard deviation of 0.8 bikes. The cost of setting up for the sidewalk sale is $200 a) Define random variables and use them to express the bicycle shops net profit. b) Whats the mean of the net profit? c) Whats the standard deviation of the net profit? d) Do you need to make any assumptions in calculating the mean? How about the standard deviation? wi
Answer #1

1. A) as the probability of cars being passing inspected is independent, the probability of my car passing inspection = 0.75 * 0.75 + 0.75 * 0.25 = 0.75

B) probability of my car not passing inspection = 1-0.75 = 0.25

C) probability that both cars pass = 0.75*0.75 = 0.5625

D) probability atleast one car passes = 0.75*0.75 + 0.75*0.25 + 0.25*0.75 = 0.9375

2) probability atleast one of them is defective = 1- probability none is defective = 1-0.98^4 = 0.0776394